OpenStudy (anonymous):

Please help FAST!! Part 1: Solve each of the quadratic equations below. Show your work. x2 − 16 = 0 and x2 = −2x + 24 Part 2: Describe what the solution(s) represent to the graph of each. Part 3: How are the graphs alike? How are they different? Thanks for any responses!

OpenStudy (anonymous):

\[x ^{2} - 16 = 0\] Can be written as : \[(x ^{2}- 4^{2}) = 0\] Now, by using the formula = > \[(a ^{2}-b ^{2}) = (a+b) (a-b)\] , we can write the equation as : \[(x+4) (x-4)\] = 0 Therefore, x =+4 or x = -4.

OpenStudy (anonymous):

Now, as for the second equation, bring all the terms to one side: \[x ^{2} + 2x - 24 = 0\] Now, by using the splitting method / quadratic formula, we get x. \[x ^{2} - 6x - 4x - 24 = 0\] \[x(x+6) - 4( x+ 6) = 0\] \[(x-4) ( x+6) = 0\] Therefore, x = 4 or x = -6.

OpenStudy (anonymous):

Thanks @Jas9420